A 7 m diameter spherical ball is known to emit radiation at a rate of 340 W when its surface temperature is 880 K. Determine the average emissivity of the ball at this temperature. 𝜎 = 0.1714 x10-8 W/m2K 4 .


1
Expert's answer
2021-07-27T12:41:01-0400

"e = \frac{P}{\sigma AT^4}"

"e = \frac{340}{(0.25\times 0.1714 \times 10^{-8} \times7^2\times880^4}=0.027"